OpenStudy (anonymous):

Verify the identity. cot x-pi/2= -tan x

OpenStudy (e.mccormick):

What have you tried on this?

OpenStudy (anonymous):

Don't I have to use the 90 degree reflection ?

OpenStudy (e.mccormick):

Sum to product would work nice.

OpenStudy (anonymous):

so help me set it up and I'll solve it because I am beyond confused haha.

OpenStudy (anonymous):

Is it cot x - pi/2 OR Cot(x - pi/2) both r quite different

OpenStudy (e.mccormick):

Ah, true.

OpenStudy (anonymous):

cot(x-pi/2)=tan x sorry about that

OpenStudy (anonymous):

Yeah so pi/2 is actually 90 degrees. U know that cot(90-x) = Tan x ?

OpenStudy (anonymous):

ok so I was right about the 90 degree ref. awesome ! haha

OpenStudy (anonymous):

haha yeah :P so if cot(90-x) = Tan x, Cot(x-90) would mean that x is more than 90 and hence x would be in quadrant 2, where Tan is negative. Hence cot(x-90) = Tan x. (Provided 90 < x < 180)

OpenStudy (anonymous):

OR u can say that cot(x-90) = -cot(90-x) = - Tan x

OpenStudy (anonymous):

(((((((: ok I finally understand ok thank you! I just need to stop second guessing myself.

OpenStudy (anonymous):

i also need help on this last one : Write the expression as the sine, cosine, or tangent of an angle. sin 9x cos x - cos 9x sin x sin 10x cos 8x sin 8x cos 10x

OpenStudy (e.mccormick):

Ah, yes. Even/Odd properties. That would certainly make it quicker!

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

Ok. @mheiges1 do you know that sin(a-b) = sin a cos b - cos a sin b ?

OpenStudy (anonymous):

Where do I find the values of A and B to plug in ?

OpenStudy (anonymous):

@mheiges1 u don't need to find. sin 9x cos x - cos 9x sin x is a form of the above identity. where sin a = sin 9 x and cos b = cos x hence a = 9x, b = x so it goes as simple as sin(9x-x) = sin 8x.

OpenStudy (anonymous):

sin 8x = sin (9x-x) sin(9x-x) = sin 9x cos x - cos 9x sin x.

OpenStudy (anonymous):

Cot(x-pi/2)=-cot(pi/2-x)=--tanx because cot(-y)=-Coty [-y is in 4th quadrant & in 4th quadrant cot is negative.

OpenStudy (anonymous):

oh ok I gotcha! thanks again dude. you rock @luckythebest

OpenStudy (anonymous):

yeah thanks